OpenStudy (anonymous):

6+2 (3j-2) = 4(l+j) please show work

OpenStudy (anonymous):

are you asked to solve for j

OpenStudy (anonymous):

In 4(1+j)or 4(l+j)

OpenStudy (anonymous):

what the question ask you to do

OpenStudy (anonymous):

@Celeste7 4(L+j)

OpenStudy (anonymous):

Solve forj

OpenStudy (aripotta):

wouldn't you have to distribute first, then add?

OpenStudy (geometry_hater):

Are you sure that L is actually L and not a 1?

OpenStudy (anonymous):

It could not be like because you will have 2 different variables l and j

OpenStudy (anonymous):

Sorry it is 1

OpenStudy (geometry_hater):

Ok just do the same thing up there just 4(l+j) would turn out to be 4+4j

OpenStudy (geometry_hater):

6+2 (3j-2) Here you actually do distribute the 2 first)......After you got all you stuff just combine like terms and then isolate j

OpenStudy (anonymous):

6+2(3j-2)=4(1+j) 6+6j-4=4+4j 6j-4+6=4+4j 6j+2=4+4j 6j-4j=4-2 2j=2 J=2/2 J=1. And you can check it by replacing the answer by the variable

OpenStudy (anonymous):



OpenStudy (anonymous):

If you need another thing you tell it ill help you

OpenStudy (anonymous):

I show you all the process so you did not confuse

OpenStudy (mayankdevnani):

any problem

OpenStudy (anonymous):

thanks

OpenStudy (mayankdevnani):

all right